OpenStudy (anonymous):

when p(x)=ax^3+bx+c is divided by (x-1) the remainder is -4. when p(x) is divided by (x^2-4) the remainder is -4x+3.find a b and c

OpenStudy (anonymous):

did you attempt to do the long division?

OpenStudy (anonymous):

you there?

OpenStudy (raden):

use remainder theorema, : p(x) = ax^3+bx+c if p(x) divided (x-1), the remainder is -4. it means p(1) = -4 a(1)^3 + b(1) + c = -4 a + b + c = -4 ... (1) next, x^2-4 = (x+2)(x-2) if p(x) divided (x+2), the remainder is -4x+3. it means p(-2) = -4(-2) + 3 p(-2) = 11 a(-2)^3 + b(-2) + c = 11 -8a - 2b + c = 11 ... (2) if p(x) divided (x-2), the remainder is -4x+3. it means p(2) = -4(2) + 3 p(2) = -5 a(2)^3 + b(2) + c = 11 8a + 2b + c = -5 ... (3) now, there are three equation with 3 variables... just find solution of them... maybe u can use subtitution and elimination method

OpenStudy (anonymous):

yeah i am here

OpenStudy (anonymous):

that's actually the long way in this case... but whatever.

OpenStudy (raden):

sorry, there is a typo error before get 3th equation :) it should be : a(2)^3 + b(2) + c = -5 (not 11) 8a + 2b + c = -5 ... (3)

OpenStudy (anonymous):

thanks so much

OpenStudy (raden):

welcome...